OpenStudy (anonymous):

someone please help Which of the following describes the non-rigid transformation in the function shown below? y-1=-(3x+1)^2 A. The graph is stretched horizontally to 1/3 the original width B. The graph is stretched vertically by a factor of 3. C. The graph is shifted 3 units down. D. The graph is reflected across the x-axis.

OpenStudy (anonymous):

What can you tell me about transformations?

OpenStudy (anonymous):

i dont really know anything about them, thats why i need help

OpenStudy (anonymous):

Ok well then lets start with the basics and work our way up.

OpenStudy (anonymous):

That work?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Sorry, Didn't see you reply to this. Are you still interested in working through transformations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We start by looking at something with no transformation at all: y = x^2 Then we compare it to a transformation: y = x^2 + 1 How will the y values differ from the first version of the function for a given value of x?

OpenStudy (anonymous):

moved to the left 1??

OpenStudy (anonymous):

Nope. If I plug in a 0 for x in the first one, I get y = 0, But if I plug in a 0 for x in the second one I get y = 1. So the graph has been shifted up 1 unit for the same value of x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And if we re-write it as: y - 1 = x^2 It's still the same thing. A parabola shifted up one unit. So now the vertex is at (0,1) instead of (0,0).

OpenStudy (anonymous):

Now if we see that x = (x-0) We can also re-write it as: y - 1 = (x-0)^2

OpenStudy (anonymous):

which you know from your other problem is an easy way to find the vertex (0,1)

OpenStudy (anonymous):

so in the problem, were looking at the -(3x) part of the problem, right?

OpenStudy (anonymous):

We will do that part now yeah.

OpenStudy (anonymous):

So it's -(3x+1)

OpenStudy (anonymous):

so wats the transformation for -(3x+1)